GUIDELINES FOR TYPESETTING A CAMERA-READY MANUSCRIPT BY COMPUTER

AUTHOR’S NAME University Department, University Name, Address City, State ZIP/Zone, Country E-mail: username@abc.def.gh

and SECOND AUTHOR’S NAME Group, Company, Address, City, State ZIP/Zone,Country

ABSTRACT

This text was produced using L^{A}TEX and formatted using pdproc.sty . Please adhere as closely to these guidelines as possible.

1. General Appearance

The typeset manuscript must be in its final form and of good appearance because it will be filmed and printed directly without any editing. It is essential that the ‘camera-ready copy’ be absolutely clean and unfolded. It should be evenly printed on a high-resolution printer (300 dots per inch or higher). There should not be corrections made on the printed pages, nor should adhesive tape cover any typeset lettering. Photocopies are not acceptable.

1.1. Text Formatting

A font with serifs, such as New Century Schoolbook, Times, or L^{A}TEX’s Computer Modern Roman font, should be used throughout. Unless otherwise specified, the font should appear in a plain style, i.e. not bold or italic.

The document’s title should be in \underline{capital letters} and in 12 point bold text, with a baselineskip (or ‘leading’) of 15 points. The name also in \underline{capital letters}, address and e-mail address of each author should be in 10 point text with a baselineskip of 13 points; the postal address should be in italics. The Abstract should have an indentation of 0.5 inches (12 mm) on the left and right margins and be in 10 point text with a baselineskip of 13 points. The main text of the article should be typeset in 12 point, preferably with a baselineskip of 15 points. The text area is 6 inches (15.2 cm) across and 8.6 inches (21.8 cm) deep, excluding page numbers. Final pagination will be done by the publisher. (Please manually adjust your page and paragraph breaks to ensure that the page length is consistent and that isolated lines of text do not occur.)

1.2. Photoreduction of Manuscript

Note that the manuscript will be printed ∼ 15% smaller than the original. Therefore please ensure that all text, including captions and labeling in figures, is large enough as to be easily legible in the printed version.

1.3. Section Headings

Section headings should be in 12 point bold, with uppercase letters at the start of major words, the remaining letters being lower case. Sub-headings are to be similarly typeset but in italics. For each section or sub-heading, allow a space of about 0.25 inches (6 mm or 17 points) above it and 0.16 inches (4 mm or 12 points) below.

2. Equations

Displayed equations should be centralized and numbered consecutively, with the equation number flush right (i.e. right-justified) and enclosed in parentheses. Equations should be referred to in the text as Eq. (X), where X is the equation number. In multiple-line equations, the number should be given on the last line. Please ensure that equations are numbered correctly, without repetition, and that no important equations are omitted from the numbering scheme.

Equations should be set in the same font size as the main text, with superscripts and subscripts 2–3 points smaller.

3. Illustrations and Photographs

Illustrations must be clear and unfolded, and their print quality even and dark enough for reproduction. It is usually sufficient that the figures be generated using modern graphics software, then laserprinted.

Please avoid mounting figures with adhesive tape. If tape is absolutely necessary then ensure that it does not cover any typeset lettering nearby. Figures are to be embedded in the text near where they are first referenced, and within the text area specified in Sec. 1.1. (Alternatively, a suitable space may be left in the text for figures to be inserted manually later.) Captions must be set below the figure, in 10 point text with a baselineskip of 13 points, and sequentially numbered with Arabic numerals ( see Fig: 1). Black and white photographs are strongly preferred and must be sharp.

Figure 1: Radiative Processes for the CP Eigenstates ^{2)}.

If you wish to ‘embed’ a postscript figure in the file, then remove from template file the % mark from the declaration of the postscript figure within the figure description and change the filename to an appropriate one. Also change the \usepackage{pdproc} into \usepackage{pdproc,epsfig} You may need to play around with this as different computer systems appear to use different commands.

Next adjust the scaling of the figure until it’s correctly positioned (sometimes using \centering helps), and remove the declarations of the lines and any anomalous spacing.

If instead you wish to use some other method, then leave the right amount of vspace in the figure declaration to accommodate your figure (remove the lines and change the space in the example) and paste the hard copy figure on to the space in the final hard copy.

Table 1: Γ(K → ππγ) for the K_{S}^{0}, K_{L}^{0} and K^{−} mesons.
Meson Γ(π^{+}π^{−}) s^{−1}  Γ(π^{+}π^{−}γ) s^{−1}

K_{S}^{0}  0.769 × 10^{10}  5.46 × 10^{7}

No DE observed, nor (IB)-E1 interference, despite large statistics, for E_{γ}^{∗} > 20MeV.

K_{L}^{0}

3.93 × 10^{4}

0.90 × 10^{3}

DE prominent, exceeding IB over the range of measurement 20 < E_{γ}^{∗} < 160MeV.

(DE = 0.62 × 10^{3})

4. Tables

Tables should be placed in the text near where they are first referenced. Captions should be placed above the tables and sequentially numbered within the text. Set captions in 10 point, with a baselineskip of 13 points (see Table:1).

5. Acknowledgements

Acknowledgements should appear just before the references.

6. References

1) J. F. Donoghue and B. R. Holstein, Phys. Rev. D25 (1982) 2015.

2) M. L. Cohen and P. W. Anderson, in Superconductivity in d- and f-Band Metals, ed. D. H. Douglas (AIP, New York, 1972).

3) H. Krebs, Fundamentals of Inorganic Crystal Chemistry (McGraw-Hill, London, 1968), p. 160.

It is preferred that references in the bibliography be cited in the text using a superscript number with a parenthese; for example, if we cite the paper by Cohen and Anderson the reference number appears like this^{2)}. All references should include initials and last name of the author(s), title of publication (in italics), volume (in bold), year of publication of paper in the journal/book, and page numbers.

7. Points to Note

a. Please ensure that quotation marks are paired correctly, e.g. “good quotes” rather than ”bad quotes”.

b. Italicized words should not be done in L^{A}TEX’s mathmode as this will result in unusual spacing of letters, e.g. compare the difference between features and features (the latter being done in mathmode).

8. Footnotes

Footnotes should be typeset in 10 point text at the bottom of the page where they are cited. Use lower-case letters for the superscripts.^{a}

aOur style file increments the letter using the usual footnote command.